A retailer spent $48 to purchase a number of special mugs. Two of them were broken in the store, but by selling each of the remaining mugs for $3 above the original cost per mug, she made a total profit of $22.
Construct an equation that will allow is to solve for the number of mugs, denoted by n, that were originally purchased.

Answer by srinivas123(23)

A retailer spent $48 to purchase a number of special mugs. Two of them were broken in the store, but by selling each of the remaining mugs for $3 above the original cost per mug, she made a total profit of $22.
Construct an equation that will allow is to solve for the number of mugs, denoted by n, that were originally purchased.
let the number of mugs be n
so cost of each mug is 48/n.
cost of mug for selling = 48/n +3.(three more)
total number of mugs sold = n-2.
total amount he got after selling mugs = 48 + 22 = 70.00
so ,
(48/n+3) (n-2) = 70
solving it we get n=12.
so the number of mugs initially bought are 12 mugs.
two mugs got damaged so total sold =10.
cost of each mug sold is 7.